Find the maximum distance the spring is compressed. Assume a spring constant of 1.0 x 106N/m.​​​​​​


1
Expert's answer
2021-06-25T10:44:36-0400

"F=kx," "\implies"

"x=\frac Fk,"

"x=\frac{F}{1\cdot 10^{6}}=10^{-6}F~(m)," where F is the force.